# Scheduling problem of aircrafts on a same runway and dual runways

### v3

@@ I Introduction
In this paper, the scheduling problems of landing and takeoff aircraft on a same runway and dual runways are addressed. In most of existing works, aircraft were considered based on the ICAO (International Civil Aviation Organization) separations, which includes $3$ classes of aircraft and there are totally $3\times 3=9$ kinds of pairwise combinations of landing and takeoff aircraft. In this paper, our work can be used to the RECAT-EU (Recovery Assistance for Cohesion and the Territories of Europe) systems, which might include $6$ or more classes of aircraft and there might be totally no smaller than $6\times 6=36$ kinds of pairwise combinations of landing and takeoff aircraft. The scheduling problems studied in this paper are much more complicated than the existing works based on ICAO separations. In contrast to the approaches based on mixed-integer optimization models in existing works, our approach focuses on the minimum separation times between aircraft by introducing some necessary assumptions and new concepts including relevance, breakpoint aircraft, path and class-monotonically-decreasing sequence. Four scheduling problems are discussed including landing scheduling problem, takeoff scheduling problem, and mixed landing and takeoff scheduling problems on a same runway and on dual runways. Two real-time optimal algorithms are proposed to find the optimal aircraft sequences to minimize the given objective function. Numerical examples are presented to show the effectiveness of the algorithms. In particular, when $100$ aircraft are considered, by using the algorithms in this paper, the optimal solution can be obtained in less than $5$ seconds, while by using the CPLEX software to solve the mix-integer optimization model, the optimal solution cannot be obtained within $1$ hour.
@@
  • The first contribution is that this paper establishes a new theoretical framework for scheduling problems of aircraft, which is completely different from the framework of mixed-integer optimization problem.
  • The second contribution is that optimal solutions can be obtained for scheduling problems of aircraft by using our proposed algorithms, which is different from existing works as well, where optimal solutions can rarely be obtained and the final optimal errors are usually unknown.
@@
  • The fourth contribution is that our work can be used for the RECAT-EU systems, which might include $6$ or more classes of aircraft. In existing works based on the ICAO separations, aircraft are usually classified into $3$ classes. The scheduling problems studied in this paper are much more complicated than the existing works based on ICAO separations.
@@ II Problem formulation
$\begin{array}[]{lll}\min F(\phi,Sr(\phi))\\ \mbox{Subject to}~{}~{}~{}S_{k}\in Tf_{k}=[f_{k}^{\min},f_{k}^{\max}]\\ \hskip 56.9055ptk=1,\cdots,n,\\ \hskip 56.9055ptS_{j}-S_{i}\geq Y_{{i}{j}},~{}\forall i<j,\\ \hskip 56.9055pti,j=1,\cdots,n,\end{array}$ (6)
@@ III Landing scheduling on a single runway
Based on the landing separation time standards at Heathrow Airport and the understanding of the physical landing process, we propose the following assumptions.
@@
The quantities $\rho_{1}$ and $\rho_{2}$ are defined to emphasize the differences of the aircraft of classes $3$ and $5$ from the aircraft of other classes.
Remark 2
Though different airports might have different landing parameters, the analysis idea in this paper might be still valid except some special requirements of the airport.
@@
Let $\Theta_{0}=\{\langle i,k,j,h\rangle\mid i<j<k,i<h<k,i,j,h,k\in\mathcal{I}\}$ be a class sequence set such that $T_{ik}+T_{0}\geq T_{ij}+T_{jk}$ for any $\langle i,k,j,h\rangle\in\Theta_{0}$ and $T_{ik}+T_{0}<T_{ij}+T_{jk}$ for any $\langle i,k,j,h\rangle\notin\Theta_{0}$. Let $\Theta_{1}=\{\langle i,k,j,h\rangle\mid k<i<j,k<h<j,i,j,h,k\in\mathcal{I}\}$ be a class sequence set such that $T_{ik}+T_{0}\geq T_{ij}+T_{jk}$ for any $\langle i,k,j,h\rangle\in\Theta_{0}$ and $T_{ik}+T_{0}<T_{ij}+T_{jk}$ for any $\langle i,k,j,h\rangle\notin\Theta_{0}$.
In Definition 6, the sets $\Theta_{0}$ and $\Theta_{1}$ are discussed for some scenarios when one breakpoint is split into two breakpoints or two breakpoints are merged into one breakpoint, and more scenarios, e.g., the split of one breakpoints into more than two breakpoints, can be analyzed in a similar way, all of which can be used to reduce the computation cost of our algorithms proposed later.
@@
Under Assumption 5, consider landing sequences $\phi_{1}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ and $\phi_{2}=\langle Tcf_{1},Tcf_{3},Tcf_{4},Tcf_{2},Tcf_{5}\rangle$, where $\langle cl_{1},cl_{2},cl_{3},cl_{4}\rangle\in\Theta_{0}$ and $cl_{5}\leq cl_{4}<cl_{2}$. Since $\langle cl_{1},cl_{2},cl_{3},cl_{4}\rangle\in\Theta_{0}$, $cl_{1}<cl_{3}<cl_{2}$, implying that $cl_{2}\geq 3$, and hence $T_{cl_{1}cl_{3}}+T_{cl_{4}cl_{2}}\leq T_{cl_{1}cl_{2}}+T_{0}=T_{cl_{1}cl_{2}}+T_{cl_{2}cl_{3}}$. Note that $T_{cl_{4}cl_{5}}\geq T_{cl_{2}cl_{5}}=T_{0}$. Therefore, $F(\phi_{1})\geq F(\phi_{2})$. If the inequality $cl_{5}\leq cl_{4}<cl_{2}$ does not hold, it is hard to determine which of $F(\phi_{1})$ and $F(\phi_{2})$ is larger without more information. From this example, it can be observed that the class sequence in $\Theta_{0}$ might be optimal in some situations but might not be in some other situations which is additionally related to the orders of other aircraft. It should be noted that when the conditions of the class sequence sets $\Theta_{0}$ and $\Theta_{1}$ do not hold, then the two breakpoints can be merged into one breakpoints to reduce the value of the objective function $F(\cdot)$.
@@
Consider two sequences $\phi_{1}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ and $\phi_{2}=\langle Tcf_{1},Tcf_{3},Tcf_{4},Tcf_{2},Tcf_{5}\rangle$, where $(cl_{1},cl_{2},cl_{3},cl_{4},cl_{5})=(i_{1},i_{2},i_{3},i_{4},i_{5})$. Suppose that Assumptions 1-5 hold for all landing aircraft sequences, and ${i_{1}}\geq{i_{2}}\geq{i_{3}},{i_{4}}\geq{i_{5}},$ and $i_{1},i_{2},i_{3},i_{4},i_{5}\in\mathcal{I}$. If $i_{2}=j=k$ for some $j\in\{i_{1},i_{3}\}$ and some $k\in\{\rho_{4},\rho_{5}\}$, then $F(\phi_{1})=F(\phi_{2})$.
@@
Proposition 3 shows that if the consecutive aircraft are of the same class and one of them needs to be moved, all the consecutive aircraft with the same class are usually needed to be move together simultaneously.

### v5

@@ I Introduction
In this paper, the scheduling problems of landing and takeoff aircraft on a same runway and dual runways are addressed. In most of existing works, aircraft were considered based on the ICAO (International Civil Aviation Organization) separations, which includes $3$ classes of aircraft and there are totally $3\times 3=9$ kinds of pairwise combinations of landing and takeoff aircraft. In this paper, our work can be used to the RECAT-EU (Recovery Assistance for Cohesion and the Territories of Europe) systems, which might include $6$ or more classes of aircraft and there might be totally no smaller than $6\times 6=36$ kinds of pairwise combinations of landing and takeoff aircraft. The scheduling problems studied in this paper are much more comprehensive and complicated than the existing works based on ICAO separations. In contrast to the approaches based on mixed-integer optimization models in existing works, our approach focuses on the minimum separation times between aircraft by introducing some necessary assumptions and new concepts including relevance, breakpoint aircraft, path and class-monotonically-decreasing sequence. Four scheduling problems are discussed including landing scheduling problem, takeoff scheduling problem, and mixed landing and takeoff scheduling problems on a same runway and on dual runways. Two real-time optimal algorithms are proposed to find the optimal aircraft sequences to minimize the given objective function. Numerical examples are presented to show the effectiveness of the algorithms. In particular, when $100$ aircraft are considered, by using the algorithms in this paper, the optimal solution can be obtained in less than $5$ seconds, while by using the CPLEX software to solve the mix-integer optimization model, the optimal solution cannot be obtained within $1$ hour.
@@
  • The first contribution is that this paper studies the interaction mechanism of the minimum separation times between aircraft in sequences and establishes a new theoretical framework for scheduling problems of aircraft, which is completely different from the framework of mixed-integer optimization problem.
  • The second contribution is that optimal solutions can be obtained for scheduling problems of aircraft by excluding most of the non-optimal solutions and narrowing down the search for the optimal solutions within a small range. This is also different from existing works, where optimal solutions can rarely be obtained and the final optimal errors are usually unknown.
@@
  • The fourth contribution is that our work can be used for the RECAT-EU systems, which might include $6$ or more classes of aircraft. In existing works based on the ICAO separations, aircraft are usually classified into $3$ classes. The scheduling problems studied in this paper are much more comprehensive and complicated than the existing works based on ICAO separations.
@@ II Problem formulation
$\begin{array}[]{lll}\min F(\phi,Sr(\phi))\\ \mbox{Subject to}~{}~{}~{}S_{k}\in Tf_{k}=[f_{k}^{\min},f_{k}^{\max}],\\ \hskip 56.9055ptk=1,\cdots,n,\\ \hskip 56.9055ptS_{j}-S_{i}\geq Y_{{i}{j}},~{}\forall i<j,\\ \hskip 56.9055pti,j=1,\cdots,n,\end{array}$ (6)
@@
In this paper, we will discuss the scheduling problems of aircraft under four typical scenarios including the landing/takeoff scheduling problem and the mixed landing and takeoff scheduling problems on a same runway and dual runways, which can be extended to the more general scheduling problems based on the analysis approach in this paper.
@@ III Landing scheduling on a single runway
Based on the minimum landing separation time standards at Heathrow Airport [30] and the understanding of the physical landing process, we propose the following assumptions.
@@
Though different airports might have different parameters, the analysis idea in this paper might be still valid except some special requirements of the airport.
@@
Proof: When $i\geq k$, from Assumptions 1 and 3, $T_{ik}\leq 1.5T_{0}<2T_{0}\leq T_{ij}+T_{jk}$. When $i\leq k\leq j$, from Assumption 2, $T_{ik}\leq T_{ij}<T_{ij}+T_{jk}$. When $j\leq i\leq k$, $T_{ik}\leq T_{jk}<T_{ij}+T_{jk}$. When $i\leq j\leq k$, from Assumption 2, $T_{ik}<T_{ij}+T_{jk}$.
@@
Proof: If $k<i-3$, $S_{ki}>3T_{0}\geq T_{cl_{k}cl_{i}}$ and hence the aircraft $Tcf_{i}$ is not relevant to $Tcf_{k}$. If $k=i-3$ or $k=i-2$, from Assumption 2, $S_{ki}=S_{k(k-1)}+S_{(k-1)i}\geq T_{cl_{k}cl_{k-1}}+T_{cl_{k-1}cl_{i}}>T_{cl_{k}cl_{i}}$. Therefore, if $k=i-3$ or $k=i-2$, the aircraft $Tcf_{i}$ is not relevant to the aircraft $Tcf_{k}$. From the statement (1), the statements (2) and (3) naturally hold.
@@
Proof: By simple calculations, this lemma can be proved and hence its proof is omitted.
@@
Proof: (1) Since $h_{1}=1$, $cl_{s_{1}}\leq 2$. If $cl_{s_{1}}>2$ and the number of the aircraft of class no larger than 2 in $\phi_{2}$ is larger than $1$, from Assumption 4, $T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}cl_{s_{1}+1}}>0.5T_{0}$. It follows from Assumptions 1 and 3 that
$\begin{array}[]{lll}F(\Phi_{a})-F(\Phi_{b})\\ =T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}\\ +T_{h_{1}cl_{i+1}}+T_{cl_{i-1}h_{1}}-T_{cl_{i-1}cl_{i+1}}\\ <-0.5T_{0}-T_{0}+1.5T_{0}<0,\end{array}$
where $T_{cl_{i-1}h_{1}}=T_{cl_{i-1}cl_{i+1}}$ since $cl_{i+1}\leq h_{1}=1$. If $cl_{s_{1}}=2$, noting that $Tcf_{s_{1}+1}\neq Tcf_{i}$, by a similar approach, it can also be obtained that $F(\Phi_{a})-F(\Phi_{b})<0$.
(2) Since $s_{2}>1$, either $cl_{i-1}$ or $cl_{i+1}$ is equal to $2$. We assume that $cl_{i+1}=2$ and the case of $cl_{i-1}=2$ can be discussed in the same way. Since $h_{1}=2$ and $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\in E$,
$\begin{array}[]{lll}F(\Phi_{a})-F(\Phi_{b})\\ =T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}+T_{h_{1}cl_{i+1}}\\ +T_{cl_{i-1}h_{1}}-T_{cl_{i-1}cl_{i+1}}\\ =T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}+T_{h_{1}cl_{i+1}}\\ =T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}+0.5T_{0},\end{array}$
where $T_{cl_{s_{1}}h_{1}}=T_{0}$ since $h_{1}<cl_{s_{1}}$ and $T_{h_{1}cl_{i+1}}=1.5T_{0}$ since $h_{1}=cl_{i+1}=2$. Note that $0.5T_{0}-\delta\leq T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}cl_{s_{1}+1}}<0.5T_{0}$ since $\langle cl_{s_{1}},cl_{s_{1}+1}\rangle\in E$. Hence, $0<F(\Phi_{a})-F(\Phi_{b})\leq\delta$.
(3)-(5) The proofs of statements (3)-(5) can be proved by a similar approach to the proofs of statements (1) and (2).
@@
Proof: (1) Since $(cl_{i},cl_{n_{1}})=(\rho_{2},\rho_{2}-1)$, from Lemma 3, under Assumptions 3 and 4, it can be checked that $F(\Phi_{a})-F(\Phi_{b})=0$.
(2) When $(cl_{n_{1}},cl_{n_{1}+1})=(\rho_{2}-1,\rho_{2})$, $cl_{i}\neq\rho_{2}$ for all $i\in\{1,2,\cdots,n_{1}-1\}$. It follows that $f_{\rho_{2}}(\Phi_{a})+f_{\rho_{2}-1}(\Phi_{a})-f_{\rho_{2}}(\Phi_{b})-f_{\rho_{2}-1}(\Phi_{b})=\delta$. Note that $f_{k}(\Phi_{a})=f_{k}(\Phi_{b})$ for all $k\neq\rho_{2}-1,\rho_{2}$. Therefore, $F(\Phi_{a})-F(\Phi_{b})>0$. When $(cl_{n_{1}},cl_{n_{1}+1})\neq(\rho_{2}-1,\rho_{2})$, from Assumptions 2 and 4, $T_{cl_{n_{1}}cl_{(n_{1}+1)}}\geq 1.5T_{0}>T_{0}+2\delta$. From Assumption 1, there are at most two classes of aircraft satisfying $T_{ii}=T_{0}+\delta$. Thus, during the transition from $\Phi_{a}$ to $\Phi_{b}$, there are at most two aircraft such that the separation times between them and their trailing aircraft become $T_{0}+\delta$ from $T_{0}$ or $T_{cl_{n_{1}}cl_{(n_{1}+1)}}$. It follows that $F(\Phi_{a})-F(\Phi_{b})\geq T_{0}+2\delta-T_{0}-2\delta>0$.
(3) Similar to the proof of statement (2), during the transition from $\Phi_{a}$ to $\Phi_{b}$, there are at most two aircraft of classes $\rho_{1}$ and $\rho_{2}$ such that the separation times between them and their trailing aircraft become $T_{0}+\delta$ from $T_{0}$. When $\phi_{2}$ has aircraft of class no larger than $2$, we have
$\begin{array}[]{lll}F(\Phi_{a})-F(\Phi_{b})\\ \geq f_{1}(\Phi_{a})+f_{2}(\Phi_{a})-f_{1}(\Phi_{b})-f_{2}(\Phi_{b})-2\delta\\ \geq T_{cl_{n_{1}}cl_{n_{1}+1}}-1.5T_{0}-2\delta.\end{array}$
When $\phi_{2}$ has no aircraft of class no larger than $2$, we have
$\begin{array}[]{lll}F(\Phi_{a})-F(\Phi_{b})\\ \geq f_{1}(\Phi_{a})+f_{2}(\Phi_{a})-f_{1}(\Phi_{b})-f_{2}(\Phi_{b})-T_{0}-2\delta\\ \geq T_{cl_{n_{1}}cl_{n_{1}+1}}-T_{0}-2\delta.\end{array}$
Note from Assumption 4 that $T_{cl_{n_{1}}cl_{n_{1}+1}}>1.5T_{0}+2\delta$ since $cl_{n_{1}}\in\{1,2\}$. Then it follows that $F(\Phi_{a})-F(\Phi_{b})>0$.
@@
Proof: (1) Calculate $F(\Phi_{a})-F(\Phi_{b})$ in three parts: the separation times between the breakpoint aircraft and their trailing aircraft, the separation times between the aircraft of classes $\rho_{1}$ and $\rho_{2}$ and their trailing aircraft, and the separation times between the last aircraft in $\Phi_{a}$ and $\Phi_{b}$ and their trailing aircraft. It can be obtained that $F(\Phi_{a})=F(\Phi_{b})-T_{d1}+T_{cl_{sc_{s}}1}-\sum_{i=1}^{s}[\mathrm{Exs}_{\rho_{1}}(\phi_{i})+\mathrm{Exs}_{\rho_{2}}(\phi_{i})]\delta+\sum_{i=1}^{s-1}[T_{cl_{ic_{i}}cl_{(i+1)1}}-T_{cl_{ic_{i}}1}]$.
(2) Since $(cl_{j},cl_{kc_{k}},cl_{(k+1)1})=(\rho_{2},\rho_{2}-1,\rho_{2})$ for some $j\in\{k1,k2,\cdots,k(c_{k}-1)\}$ and all $k\in\{1,2,\cdots,s-1\}$, the subsequence $\phi_{1}$ contains at most the aircraft of classes $\rho_{2}-1$, $\rho_{2}$, $\cdots,\eta$, each of the subsequences $\phi_{2},\phi_{3},\cdots,\phi_{s-1}$ contains the aircraft of classes $\rho_{2}-1$ and $\rho_{2}$, and the subsequence $\phi_{s}$ contains at most the aircraft of classes $\rho_{2}$, $\rho_{2}-1$, $\cdots,1$. Note that $T_{d1}-T_{cl_{sc_{s}}1}\geq 0$ since $d\leq cl_{i}$ for all $Tcf_{i}$ in $\Phi_{b}$. Thus, it can easily be checked that $F(\Phi_{a})-F(\Phi_{b})=0$ based on the statement (1).
(3) When $(cl_{ic_{i}},cl_{(i+1)1})=(\rho_{2}-1,\rho_{2})$ for some $i\in\{1,\cdots,s-1\}$, under the condition given in this statement, it follows that $T_{cl_{ic_{i}}cl_{(i+1)1}}-T_{cl_{ic_{i}}1}=\delta$ and $\mathrm{Exs}_{\rho_{2}}(\phi_{i})=\mathrm{Exs}_{\rho_{1}}(\phi_{i})=0$. When $(cl_{ic_{i}},cl_{(i+1)1})\neq(\rho_{2}-1,\rho_{2})$ for some $i\in\{1,2,\cdots,s-1\}$, from Assumptions 3, 2 and 4, if $cl_{ic_{i}}=1$, then $T_{cl_{ic_{i}}cl_{(i+1)1}}-T_{cl_{ic_{i}}1}>0.5T_{0}$, if $cl_{ic_{i}}=2$, then $T_{cl_{ic_{i}}cl_{(i+1)1}}-T_{cl_{ic_{i}}1}>2\delta$, and if $cl_{ic_{i}}\geq 3$, then $T_{cl_{ic_{i}}cl_{(i+1)1}}-T_{cl_{ic_{i}}1}\geq 0.5T_{0}$. Note that $T_{d1}-T_{cl_{sc_{s}}1}\geq 0$ since $d\leq cl_{i}$ for all $Tcf_{i}$ in $\Phi_{b}$. Based on the statement (1), it can be obtained that $F(\Phi_{a})-F(\Phi_{b})>0$.
@@
Proof: This theorem can be regarded as a special case of Theorem 2.
@@
Let $\Theta_{0}=\{\langle i,k,j,h\rangle\mid i<j<k,i<h<k,i,j,h,k\in\mathcal{I}\}$ be a class sequence set such that $T_{ik}+T_{0}\geq T_{ij}+T_{jk}$ for any $\langle i,k,j,h\rangle\in\Theta_{0}$ and $T_{ik}+T_{0}<T_{ij}+T_{jk}$ for any $\langle i,k,j,h\rangle\notin\Theta_{0}$. Let $\Theta_{1}=\{\langle i,k,j,h\rangle\mid k<i<j,k<h<j,i,j,h,k\in\mathcal{I}\}$ be a class sequence set such that $T_{ik}+T_{0}\geq T_{ij}+T_{jk}$ for any $\langle i,k,j,h\rangle\in\Theta_{1}$ and $T_{ik}+T_{0}<T_{ij}+T_{jk}$ for any $\langle i,k,j,h\rangle\notin\Theta_{1}$.
In Definition 6, the sets $\Theta_{0}$ and $\Theta_{1}$ are discussed for some scenarios when one breakpoint is split into two breakpoints or two breakpoints are merged into one breakpoint, and more scenarios, e.g., the split of one breakpoint into more than two breakpoints, can be analyzed in a similar way. All of these can be used to reduce the computation cost of our algorithms proposed later.
@@
Under Assumption 5, consider landing sequences $\phi_{1}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ and $\phi_{2}=\langle Tcf_{1},Tcf_{3},Tcf_{4},Tcf_{2},Tcf_{5}\rangle$, where $\langle cl_{1},cl_{2},cl_{3},cl_{4}\rangle\in\Theta_{0}$ and $cl_{5}\leq cl_{4}<cl_{2}$. Since $\langle cl_{1},cl_{2},cl_{3},cl_{4}\rangle\in\Theta_{0}$, $cl_{1}<cl_{3}<cl_{2}$, implying that $cl_{2}\geq 3$, and hence $T_{cl_{1}cl_{3}}+T_{cl_{4}cl_{2}}\leq T_{cl_{1}cl_{2}}+T_{0}=T_{cl_{1}cl_{2}}+T_{cl_{2}cl_{3}}$. Note that $T_{cl_{4}cl_{5}}\geq T_{cl_{2}cl_{5}}=T_{0}$. Therefore, $F(\phi_{1})\geq F(\phi_{2})$. If the inequality $cl_{5}\leq cl_{4}<cl_{2}$ does not hold, it is hard to determine which of $F(\phi_{1})$ and $F(\phi_{2})$ is larger without more information. From this example, it can be observed that the class sequence in $\Theta_{0}$ might be optimal in some situations but might not be in some other situations which is additionally related to the orders of other aircraft. It should be noted that when the conditions of the class sequence sets $\Theta_{0}$ and $\Theta_{1}$ do not hold, then the two breakpoints can be merged into one breakpoint to reduce the value of the objective function $F(\cdot)$.
@@
Proof: Using the previous lemmas and theorems, this proposition can be proved and its proof is omitted.
It should be noted that the transitions in Proposition 1 (1)(2) can be regarded as reciprocal inverse transformations of each other.
@@
Consider two sequences $\phi_{1}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ and $\phi_{2}=\langle Tcf_{1},Tcf_{3},Tcf_{4},Tcf_{2},Tcf_{5}\rangle$, where $(cl_{1},cl_{2},cl_{3},cl_{4},cl_{5})=(i_{1},i_{2},i_{3},i_{4},i_{5})$. Suppose that Assumptions 1-5 hold for all landing aircraft sequences, ${i_{1}}\geq{i_{2}}\geq{i_{3}},{i_{4}}\geq{i_{5}},$ and $i_{1},i_{2},i_{3},i_{4},i_{5}\in\mathcal{I}$. If $i_{2}=j=k$ for some $j\in\{i_{1},i_{3}\}$ and some $k\in\{\rho_{4},\rho_{5}\}$, then $F(\phi_{1})=F(\phi_{2})$.
@@
Proof: Calculating $F(\Phi_{2})-F(\Phi_{1})$, we have $F(\Phi_{2})-F(\Phi_{1})=T_{k_{1}k_{3}}+T_{k_{3}k_{2}}+(m_{3}-1)T_{k_{3}k_{3}}-T_{k_{1}k_{2}}-m_{3}T_{k_{3}k_{3}}=T_{k_{1}k_{3}}+T_{k_{3}k_{2}}-T_{k_{3}k_{3}}-T_{k_{1}k_{2}}$, where $m_{3}$ denotes the number of aircraft in $\phi_{3}$. Note that $k_{2}<k_{1}<k_{3}$. It follows that $k_{3}\geq 3$ and $k_{3}-k_{2}\geq 2$. From Assumption 4, when $k_{1}\leq 2$, $T_{k_{1}k_{3}}>1.5T_{0}+2\delta=T_{k_{1}k_{2}}+2\delta$, $T_{k_{3}k_{2}}=T_{0}$, and $T_{k_{3}k_{3}}\leq T_{0}+\delta$. Hence, $F(\Phi_{1})<F(\Phi_{2})$ when $k_{1}\leq 2$. When $k_{1}\geq 3$, it follows that $T_{k_{1}k_{3}}\geq 1.5T_{0}$ since $(k_{1},k_{3})\neq(\rho_{2}-1,\rho_{2})$, $T_{k_{3}k_{2}}=T_{k_{1}k_{2}}=T_{0}$, and $T_{k_{3}k_{3}}\leq T_{0}+\delta$. Therefore, $F(\Phi_{1})<F(\Phi_{2})$ when $k_{1}\geq 3$. This proposition is proved.
@@
Proposition 3 shows that if the consecutive aircraft are of the same class and one of them needs to be moved, all the consecutive aircraft with the same class are usually needed to be moved together simultaneously.